Edge dolocator assignment Locator-assignment-take Use the References to access important values if needed for this qontion A student weighs out a 8.07 g sample of CuCl, transfers it to a 250. ml. volumetric flask, adds enough water to dissohve it and then adds water to the 250. mL. tick mark What is the molarity of copper(II) chloride in the resulting solution? Molarnty Submit Answer Retry Entire Group 2 more group attempts remaining
media%2F346%2F346e47e0-c3d0-475d-a179-27
media%2F1e7%2F1e709a27-b55d-41c7-8867-8f
media%2F8e3%2F8e3b3321-cd06-4920-a8f7-02
Answer #1

1)

Molar mass of CUCl2,

MM = 1*MM(C) + 1*MM(U) + 2*MM(Cl)

= 1*12.01 + 1*238.0 + 2*35.45

= 320.91 g/mol

mass(CUCl2)= 8.07 g

use:

number of mol of CUCl2,

n = mass of CUCl2/molar mass of CUCl2

=(8.07 g)/(3.209*10^2 g/mol)

= 2.515*10^-2 mol

volume , V = 2.5*10^2 mL

= 0.25 L

use:

Molarity,

M = number of mol / volume in L

= 2.515*10^-2/0.25

= 0.1006 M

Answer: 0.101 M

2)

W/W% = weight of FeBr3 * 100 / total mass

= 8.10 * 100 / (8.10 + 17.5)

= 31.6 %

Answer: 31.6 %

Only 1 question at a time please